# Listen to Look into the Future: Audio-Visual Egocentric Gaze Anticipation

### 07 Dec 2023 v2

@@ Abstract
Egocentric gaze anticipation serves as a key building block for the emerging capability of Augmented Reality. Notably, gaze behavior is driven by both visual cues and audio signals during daily activities. Motivated by this observation, we introduce the first model that leverages both the video and audio modalities for egocentric gaze anticipation. Specifically, we propose a Contrastive Spatial-Temporal Separable (CSTS) fusion approach that adopts two modules to separately capture audio-visual correlations in spatial and temporal dimensions, and applies a contrastive loss on the re-weighted audio-visual features from fusion modules for representation learning. We conduct extensive ablation studies and thorough analysis using two egocentric video datasets: Ego4D and Aria, to validate our model design. We also demonstrate our model outperforms prior state-of-the-art methods by at least +1.9% and +1.6%. Moreover, we provide visualizations to show the gaze anticipation results and provide additional insights into audio-visual representation learning.
@@ 1 Introduction
A person’s eye movements during their daily activities are reflective of their intentions and goals (see [19] for a representative cognitive science study). The ability to predict the future gaze targets of the camera-wearer from egocentric videos, known as egocentric gaze anticipation, is therefore a key step towards understanding and modeling cognitive processes and decision making. Furthermore, this capability could enable new applications in Augmented Reality and Wearable Computing, especially in social scenarios - for example, providing memory aids for patients with cognitive impairments, or reducing the latency of content delivery in such AR systems. However, forecasting the gaze fixations of a camera-wearer using only the egocentric view (i.e. without eye tracking at testing time) is very challenging due to the complexity of egocentric scene content and the dynamic nature of gaze behaviors.
@@
We argue that audio signals can serve as an important auxiliary cue for egocentric gaze forecasting. Consider the example in Fig. 1. During the input video sequence, the camera view shifts from the paper held by the camera wearer to the standing speaker, who asks a question. In the anticipation sequence, the sitting speaker on the far right begins to answer the question and the camera wearer’s gaze shifts towards the sitting person’s head. The influence of audio signals on eye movements is also evidenced by neuroscience research [50]. Therefore, we address the problem of forecasting the gaze fixation of the camera-wearer in unseen future frames using a short egocentric video clip and corresponding audio. As shown in Fig. 1, the model’s ability to fuse the audio and visual cues enables it to correctly predict the future attention to the seated subject.
Though several works have addressed egocentric gaze estimation [24, 26, 25, 34, 35, 54, 33], the problem of egocentric gaze anticipation is largely understudied [65]. Moreover, no prior works on egocentric gaze modeling have explored how to leverage the audio modality. Intuitively, in the spatial dimension, the visual region (e.g. sound source) that has a stronger correlation with audio signals is more likely to be the potential future gaze target. In the temporal dimension, events in the audio signal may drive both egocentric visual capture (via head movement) and gaze movements as the camera wearer responds to new sounds. Prior approaches for audio-visual representation learning  [58, 62, 6, 21, 47, 5, 7] directly feed visual and audio embeddings into one multimodal fusion layer. However, such a fusion mechanism is not ideal under the egocentric setting, where the audio signal drives the embodied egocentric visual perception, and thus causes drastic scene change. Therefore, predicting the future gaze behavior demands a model can reason the spatial and temporal relationships between the audio and visual signals in separate.
@@
To this end, we propose a novel Contrastive Spatial-Temporal Separable (CSTS) audio-visual fusion method for egocentric gaze anticipation. Specifically, we input the egocentric video frames and the corresponding audio spectrograms into a video encoder and an audio encoder respectively. Then we leverage self-attention to develop a spatial fusion module and a temporal fusion module in parallel for modeling the spatial and temporal audio-visual correlation separately, strategically highlighting important audio-visual correlations. The output representations from the two branches are merged by channel-wise reweighting and fed into a visual decoder to predict the future gaze target. We use a multi-modal contrastive loss [2] on the reweighted representations from the fusion modules to facilitate audio-visual correspondence learning. We demonstrate the benefits of our approach on two egocentric video datasets that capture social scenarios and everyday activities: Ego4D [17] and Aria [29]. Our proposed model achieves state-of-the-art performance on egocentric gaze anticipation. Our contributions are summarized as follows:
@@
  • We introduce the first model that utilizes visual and audio signals for modeling egocentric gaze behaviors.
@@
  • We propose a novel CSTS model that leverages a spatio-temporal separable fusion module and a post-fusion contrastive learning schema to facilitate audio-visual representation learning for egocentric gaze anticipation.
@@
  • We present comprehensive experiment results on the Ego4D [17] and Aria [29] datasets. Our ablation studies show audio modality can improve the performance by +2.5% and +2.4% in F1 score on Ego4D and Aria. We then show our model also outperforms prior works by at least +1.9% and +1.6% in F1 score on the two datasets.
[…]
@@ 2 Related Work
Our work relates to three main streams of research: egocentric gaze modeling, audio-visual saliency prediction, and contrastive audio-visual representation learning. It is worth noting that no existing egocentric gaze modeling studies leverage audio signals and previous methods on audio-visual saliency prediction problem are not appropriate for egocentric gaze anticipation as detailed below.
@@
Egocentric Gaze Modeling. Modeling human gaze behavior in egocentric videos is an important topic in egocentric vision. Most prior efforts target egocentric gaze estimation [35, 33, 24, 26, 34, 25]. Huang et al. [24] propose learning temporal attention transitions from video features that reflect drastic gaze movements. Li et al. [35] and Huang et al. [25] utilize the correlation of gaze behaviors and actions, modeling them jointly with a convolutional network. Lai et al. [33] encode global scene context into a single global token and explicitly model the global-local correlations in the visual embedding for gaze estimation. In contrast, egocentric gaze anticipation, which seeks to predict future gaze targets from past video frames, addresses an understudied dimension of modeling gaze. Zhang et al. [65] introduce this task and utilize a convolutional network and a discriminator to generate future video frames, which are further used to anticipate future gaze targets. They enhance their model by adding an additional branch for gaze forecasting [66]. All previous efforts on both egocentric gaze estimation and anticipation model gaze behavior from only the visual properties of the video stream, and do not consider the relationship between audio signals and gaze behavior. In this work, we introduce the first model that leverages both visual and audio signals for egocentric gaze anticipation.
@@
Audio-Visual Saliency Prediction. Audio-visual saliency prediction is a well-studied problem in computer vision [49, 51, 48, 53, 9, 41]. Another related research topic is sound source localization [5, 52, 20, 21, 22, 23] which localizes sound source in the image/video corresponding to a given audio stream. Here, we mainly discuss previous approaches for fusing audio and visual representations. Early CNN-based approaches adopt a late-fusion strategy [57, 58, 42, 59, 60] for saliency prediction. Recently, new findings suggest audio-visual fusion at the intermediate features is a more effective way to leverage advantages of both modalities [1, 63, 8, 55] for saliency prediction. Jain et al. [27] investigate two fusion methods at the middle level which achieve new state of the art on multiple datasets. Yao et al. [64] propose to incorporate the audio signal at multiple decoder layers by using an inner-product operation. Similarly, Chang et al. [6] and Xiong et al. [62] merge audio features into visual features at multiple levels of the visual encoder. Notably, transformer-based fusion methods have not been applied in the audio-visual saliency prediction problem. Moreover, comparing to these prior works, the complex camera motion, drastic gaze movement, and the forecasting setting of egocentric gaze anticipation problem require a model that can explicitly understand how audio signals correlate with visual signals in both spatial and temporal dimensions. We implement the audio-visual fusion strategy widely used by prior work in our backbone. Experiment results in Sec. 4.2.2 demonstrate that the spatial-temporal separable fusion mechanism in our model is a more effective strategy.
@@ 3 Method
Notably, video and audio signals have correlations in both spatial and temporal dimensions: Spatially, the audio signal will have the strongest correlation with the regions of the video frames that correspond to the sound source. Temporally, in egocentric video, audio stimuli in the scene drive the wearer’s head movement and how they orient the camera’s field of view. Our key insight is thus to separate the connection between audio and visual signals into spatial and temporal correlations.
Fig. 2 demonstrates the overview of our model. We exploit the transformer-based backbone encoders $\phi(x)$ and $\psi(a)$ to extract the representations of the video frames $x$ and audio signals $a$. We then employ a Contrastive Spatial-Temporal Separable (CSTS) audio-visual fusion approach. Specifically, a spatial fusion module captures the correlation between audio embeddings and spatial appearance-based features; a temporal fusion module captures the temporal correlation between the visual and audio embeddings; and a contrastive loss is applied on fused audio-visual embeddings to facilitate the representation learning. Finally, spatially and temporally fused audio-visual features are merged and fed into a decoder for future gaze anticipation.
@@ 3.1 Audio and Visual Feature Embedding
Audio Feature Embedding. We follow [31] to adopt a sliding window approach for audio signal preprocessing. Specifically, for a video frame at time step $t_{i}$, the corresponding audio segment has a range of $[t_{i}-\frac{1}{2}\Delta t_{w},t_{i}+\frac{1}{2}\Delta t_{w}]$. We then use STFT to convert all audio segments into log-spectrograms and feed the processed audio segments into a transformer-based audio encoder $\psi(a)$. Since the audio stream has more sparse information than video stream, we follow [14, 10] to adopt a light-weighted transformer architecture for the audio encoder $\psi(a)$. In this way, $\psi(a)$ extracts $T\times M$ tokens with feature dimension $D$ from the audio inputs $a$.
@@ 3.2 Spatial-Temporal Separable Fusion
Spatial Audio-Visual Fusion. The Spatial Fusion branch identifies correlations between the audio signal corresponding to a video frame and its visual content in space. We first use convolutional operations to generate the spatial fusion audio embedding $z_{a,s}$ with dimensions $T\times 1\times D$ from the audio embedding $\psi(a)$. This allows the model to extract a holistic audio embedding within each sliding window. We then input the visual embedding $\phi(x)$ and pooled audio embedding $z_{a,s}$ into an in-frame self-attention layer $\sigma$. In this layer, we masked out all cross-frame connections and only calculate the correlations among visual tokens within each frame and the corresponding single audio token.
Therefore, the input to the spatial fusion consists of $T$ groups of visual tokens, and $T$ single audio embeddings. Formally, we have:
$\small\phi(x)=\left[\phi(x)^{(1)},...,\phi(x)^{(T)}\right],$ (1)
$\small z_{a,s}=\left[z_{a,s}^{(1)},...,z_{a,s}^{(T)}\right],$ (2)
@@
$\small z_{s}^{(i)}=\left[\phi^{(i)}(x),z_{a,s}^{(i)}\right]\in\mathbb{R}^{1\times(N+1)\times D}$ (3)
@@
$\small\sigma(z_{s}^{(i)})=Softmax\left(\bm{Q}_{s}^{(i)}{\bm{K}_{s}^{(i)}}^{T}/\sqrt{D}\right)\bm{V}_{s}^{(i)}\in\mathbb{R}^{1\times(N+1)\times D},$ (4)
@@
$\small\sigma(z_{s})=\left[\sigma(z_{s}^{(i)}),...,\sigma(z_{s}^{(T)})\right]\in\mathbb{R}^{T\times(N+1)\times D}.$ (5)
In practice, we input all tokens to the in-frame self-attention layer simultaneously, mask out cross-frame correlations and calculate Eq. 4 in one shot to speed up training. We further add two linear layers after the self-attention outputs $\sigma(z_{s})$, following the standard self-attention layer. The output of the spatial module is finally denoted as $u_{s}\in\mathbb{R}^{T\times(N+1)\times D}$.
Temporal Audio-Visual Fusion. The Temporal Fusion branch models relationships between audio and visual content across time. We apply two convolutional layers to integrate the embedding from each modality at each time step into a single token. The resulting visual and audio tokens are denoted as $z_{v,t}\in\mathbb{R}^{T\times 1\times D}$ and $z_{a,t}\in\mathbb{R}^{T\times 1\times D}$, respectively. Then we feed $z_{t}=\left[z_{v,t},z_{a,t}\right]\in\mathbb{R}^{2T\times 1\times D}$ into a cross-frame self-attention layer $\pi$:
$\small\pi(z_{t})=Softmax\left(\bm{Q}_{t}\bm{K}_{t}^{T}/\sqrt{D}\right)\bm{V}_{t}\in\mathbb{R}^{2T\times 1\times D},$ (6)
@@
Merging of Two Fusion Modules. After obtaining audio-visual representations from the two fusion modules, we merge the two branches by reweighting the output from spatial fusion with the temporal weights from temporal fusion in each channel to obtain a new representation for each modality that has been refined by multimodal spatial and temporal correlation. Specifically, we break down the output from spatial fusion $u_{s}\in\mathbb{R}^{T\times(N+1)\times D}$ into $u_{v,s}\in\mathbb{R}^{T\times N\times D}$ and $u_{a,s}\in\mathbb{R}^{T\times 1\times D}$, and the output from temporal fusion $u_{t}\in\mathbb{R}^{2T\times 1\times D}$ into $u_{v,t}\in\mathbb{R}^{T\times 1\times D}$ and $u_{a,t}\in\mathbb{R}^{T\times 1\times D}$. The reweighted visual representation is formulated as
$\small u_{v}=u_{v,s}\otimes u_{v,t}\in\mathbb{R}^{T\times N\times D},$ (7)

### 22 Mar 2024 v3

@@ Abstract
Egocentric gaze anticipation serves as a key building block for the emerging capability of Augmented Reality. Notably, gaze behavior is driven by both visual cues and audio signals during daily activities. Motivated by this observation, we introduce the first model that leverages both the video and audio modalities for egocentric gaze anticipation. Specifically, we propose a Contrastive Spatial-Temporal Separable (CSTS) fusion approach that adopts two modules to separately capture audio-visual correlations in spatial and temporal dimensions, and applies a contrastive loss on the re-weighted audio-visual features from fusion modules for representation learning. We conduct extensive ablation studies and thorough analysis using two egocentric video datasets: Ego4D and Aria, to validate our model design. We demonstrate the audio improves the performance by +2.5% and +2.4% on the two datasets. Our model also outperforms the prior state-of-the-art methods by at least +1.9% and +1.6%. Moreover, we provide visualizations to show the gaze anticipation results and provide additional insights into audio-visual representation learning. The code and data split are available on our website (https://bolinlai.github.io/CSTS-EgoGazeAnticipation/).
Keywords:
Egocentric Vision Gaze Behavior Audio-Visual Learning
footnotetext: * Equal corresponding author.
@@ 1 Introduction
A person’s eye movements during their daily activities are reflective of their intentions and goals (see [22] for a representative cognitive science study). The ability to predict the future gaze targets of the camera-wearer from egocentric videos, known as egocentric gaze anticipation, is therefore a key step towards understanding and modeling cognitive processes and decision making. Furthermore, this capability could enable new applications in Augmented Reality and Wearable Computing, especially in social scenarios – for example, providing memory aids for patients with cognitive impairments, or reducing the latency of content delivery in such AR systems. However, forecasting the gaze fixations of a camera-wearer using only the egocentric view (i.e., without eye tracking at testing time) is very challenging due to the complexity of egocentric scene content and the dynamic nature of gaze behaviors.
@@
We argue that audio signals can serve as an important auxiliary cue for egocentric gaze forecasting. Consider the example in Fig. 1. In the input sequence, the camera view shifts from the paper held by the camera wearer to the standing speaker who asks a question. Then the sitting speaker on the far right answers the question, which is captured by the audio stream. In the anticipation sequence, the camera wearer’s gaze shifts towards the sitting person’s head after hearing her response. In this case, the audio stream (the sitting person’s response) is an important stimulus that triggers this gaze movement. The influence of audio signals on eye movements is also evidenced by neuroscience research [57]. Therefore, we address the problem of forecasting the gaze fixation of the camera-wearer in unseen future frames using a short egocentric video clip and corresponding audio. As shown in Fig. 1, the model’s ability to fuse the audio and visual cues enables it to correctly predict the future attention to the seated subject.
Though many works have addressed egocentric gaze estimation [29, 30, 28, 37, 38, 63, 36], the egocentric gaze anticipation task is largely understudied [77]. Moreover, how to leverage both the visual modality and the audio modality for egocentric gaze modeling has not been explored yet.Existing methods on audio-visual learning commonly [67, 73, 6, 24, 54, 5, 7] fuse visual and audio embeddings simultaneously in time and space. However, such a fusion mechanism is not ideal under the egocentric setting, where the camera wearer’s reaction to the audio stimuli causes a drastic change of camera viewpoint. In Fig. 1, as a reaction to the question and answer, the camera wearer shifts the attention from the paper to the standing person and then to the sitting person. The viewpoint and scene also have changed because of head movement (see the first and last frame). Moreover, due to the natural delay of reaction time, the audio stimulus and gaze reaction will not occur at the same time. Therefore, predicting the future gaze behavior demands a model can (1) learn possible viewpoint and scene change driven by the audio stream over time and (2) locate the potential future gaze target in the visual space. Fusing two modalities in time and space simultaneously may result in limited performance in the two targets because of spurious audio-visual correlations. Hence, a spatial-temporal separable fusion model is a better solution for egocentric gaze anticipation task.
@@
To address the challenges in our task, we propose a novel Contrastive Spatial-Temporal Separable (CSTS) audio-visual fusion method for egocentric gaze anticipation. Specifically, we input the egocentric video frames and the corresponding audio spectrograms into a video encoder and an audio encoder respectively. Then we develop a spatial fusion module and a temporal fusion module in parallel based on self-attention mechanism for modeling the spatial and temporal audio-visual correlation separately, exactly addressing the aforementioned demands. The output representations from the two branches are merged by channel-wise reweighting and fed into a visual decoder to predict the future gaze target. We also propose a novel strategy that uses a multi-modal contrastive loss [2] on the reweighted representations (referred to as post-fusion contrastive loss) from the fusion modules to facilitate audio-visual correspondence learning. We demonstrate the benefits of our approach on two egocentric video datasets that capture social scenarios and everyday activities: Ego4D [20] and Aria [43]. The proposed model achieves state-of-the-art gaze anticipation performance on both datasets. Our contributions are summarized as follows:
  • $\bullet$
@@
    We introduce the first approach that utilizes visual and audio signals for modeling egocentric gaze behaviors.
  • $\bullet$
@@
    We propose a novel CSTS model that leverages a spatio-temporal separable fusion module and a post-fusion contrastive learning scheme to facilitate audio-visual representation learning for egocentric gaze anticipation.
  • $\bullet$
@@
    We present comprehensive experiment results on the Ego4D [20] and Aria [43] datasets. Our ablation studies show audio modality can improve the performance by +2.5% and +2.4% respectively in F1 score on Ego4D and Aria. The experiments also demonstrate our model outperforms prior state-of-the-art method by +1.9% and +1.6% in F1 score on the two datasets.
@@ 2 Related Work
Egocentric Gaze Modeling. Modeling human gaze behavior in egocentric videos is an important topic in egocentric vision. Most prior efforts target at egocentric gaze estimation [38, 36, 29, 30, 37, 28]. Huang et al. [29] propose learning temporal attention transitions from video features that reflect drastic gaze movements. Li et al. [38] and Huang et al. [28] utilize the correlation of gaze behaviors and actions, modeling them jointly with a convolutional network. Lai et al. [36] encode global scene context into a single global token and explicitly model the global-local correlations in the visual embedding for gaze estimation. In contrast, egocentric gaze anticipation, which seeks to predict future gaze targets from past video frames, addresses an understudied dimension of modeling gaze. Zhang et al. [77] introduce this task and utilize a convolutional network and a discriminator to generate future video frames, which are further used to anticipate future gaze targets. They enhance their model by adding an additional branch for gaze forecasting [76]. All previous efforts on both egocentric gaze estimation and anticipation model gaze behavior from only the visual properties of the video stream, and do not consider the relationship between audio signals and gaze behavior. In this work, we introduce the first model that leverages both visual and audio signals for egocentric gaze anticipation task.
@@
Audio-Visual Saliency Prediction. Audio-visual saliency prediction is a well-studied problem in computer vision [56, 58, 55, 61, 10, 46]. Another related research topic is sound source localization [5, 60, 23, 24, 25, 26] which localizes sound source in the image/video corresponding to a given audio stream. Here, we mainly discuss previous approaches for fusing audio and visual representations in saliency prediction problem. Early CNN-based approaches adopt a late-fusion strategy [66, 67, 47, 68, 69] for saliency prediction. Recently, new findings suggest audio-visual fusion at the intermediate features is a more effective way to leverage advantages of both modalities [1, 74, 8, 64] for saliency prediction. Jain et al. [31] investigate two fusion methods at the middle level which achieve new state of the art on multiple datasets. Yao et al. [75] propose to incorporate the audio signal at multiple decoder layers by using an inner-product operation. Similarly, Chang et al. [6] and Xiong et al. [73] merge audio features into visual features at multiple levels of the visual encoder. Notably, our problem differs from the audio-visual saliency prediction in two aspects: First, the goal of our task is forecasting gaze behavior in the future, while saliency prediction focuses more on studying human’s attention mechanism in the current video frame. Second, our problem focuses egocentric videos that capture the changing viewpoint when people respond to audio and visual stimuli, while saliency prediction uses videos captured from a fixed viewpoint, and fail to reflect gaze reaction to real-time events. Apart from the difference on problem settings, we also want to emphasize that the transformer-based fusion methods have not been applied in the audio-visual saliency prediction problem. Moreover, we propose a well-motivated spatio-temporal separable fusion module to address this challenging problem
@@ 3 Method
Notably, visual and audio signals have correlations in both spatial and temporal dimensions for gaze modeling. Spatially, the visual region (e.g., sound source) that has a stronger correlation with audio signals is more likely to be the potential future gaze target. Temporally, events in the audio signal may drive both egocentric viewpoint change (via head movement) and gaze movements as the camera wearer responds to new sounds. Our key insight is thus separating the connection of audio and visual signals into spatial and temporal correlations.
Figure 2: Overview of the proposed model. The video embeddings $\phi(x)$ and audio embeddings $\psi(a)$ are obtained by two transformer-based encoders. We then model the correlations of visual and audio embeddings using two separate branches – (1) spatial fusion, which learns the spatial co-occurence of audio signals and visual objects in each frame, and (2) temporal fusion, which captures the temporal correlations and possible gaze movement. A contrastive loss is adopted to facilitate audio-visual representation learning. We input fused embeddings into a decoder for final gaze anticipation results.
Fig. 2 demonstrates the overview of our model. We exploit the transformer-based encoders $\phi(x)$ and $\psi(a)$ to extract the representations of the video frames $x$ and audio signals $a$. We then employ a Contrastive Spatial-Temporal Separable (CSTS) audio-visual fusion approach. Specifically, a spatial fusion module captures the correlation between audio embeddings and spatial appearance-based features; a temporal fusion module captures the temporal correlation between the visual and audio embeddings; and a contrastive loss is applied on fused audio-visual embeddings to facilitate the representation learning. Finally, spatially and temporally fused audio-visual features are merged and fed into a decoder for future gaze anticipation.
@@ 3.1 Audio and Visual Feature Embedding
Audio Feature Embedding. We follow [34] to adopt a sliding window approach for audio signal preprocessing. Specifically, for a video frame at time step $t_{i}$, the corresponding audio segment has a range of $[t_{i}-\frac{1}{2}\Delta t_{w},t_{i}+\frac{1}{2}\Delta t_{w}]$. We then use STFT to convert all audio segments into log-spectrograms and feed the processed audio segments into a transformer-based audio encoder $\psi(a)$. Since the audio stream has more sparse information than video stream, we adopt a light-weighted transformer architecture (inspired by [17, 12]) for the audio encoder $\psi(a)$. In this way, $\psi(a)$ extracts $T\times M$ tokens with feature dimension $D$ from the audio inputs $a$.
@@ 3.2 Spatial-Temporal Separable Fusion
Spatial Audio-Visual Fusion. The spatial fusion branch identifies correlations between the audio signal corresponding to a video frame and its visual content in space. We first use convolutional operations to generate the audio representation $z_{a,s}$ for spatial fusion with dimensions $T\times 1\times D$ from the audio embedding $\psi(a)$. This allows the model to extract a holistic audio embedding within each sliding window. We then input the visual embedding $\phi(x)$ and pooled audio embedding $z_{a,s}$ into an in-frame self-attention layer $\sigma$. In this layer, we masked out all cross-frame connections and only calculate the correlations among visual tokens within each frame and the corresponding single audio token. Therefore, the input to the spatial fusion consists of $T$ groups of visual tokens, and $T$ single audio embeddings. Formally, we have:
$\phi(x)=\left[\phi(x)^{(1)},...,\phi(x)^{(T)}\right],$ (1)
$\vspace{-0.1cm}z_{a,s}=\left[z_{a,s}^{(1)},...,z_{a,s}^{(T)}\right],$ (2)
@@
$z_{s}^{(i)}=\left[\phi^{(i)}(x),z_{a,s}^{(i)}\right]\in\mathbb{R}^{1\times(N+1)\times D}$ (3)
@@
$\sigma(z_{s}^{(i)})=Softmax\left(\bm{Q}_{s}^{(i)}{\bm{K}_{s}^{(i)}}^{T}/\sqrt{D}\right)\bm{V}_{s}^{(i)}\in\mathbb{R}^{1\times(N+1)\times D},$ (4)
@@
$\sigma(z_{s})=\left[\sigma(z_{s}^{(i)}),...,\sigma(z_{s}^{(T)})\right]\in\mathbb{R}^{T\times(N+1)\times D}.$ (5)
In practice, we input all tokens to the in-frame self-attention layer simultaneously, mask out cross-frame correlations and calculate Eq. 4 in one shot to speed up training. We further add two linear layers after the self-attention outputs $\sigma(z_{s})$, following the standard self-attention layer design. The output of the spatial module is finally denoted as $u_{s}\in\mathbb{R}^{T\times(N+1)\times D}$.
Temporal Audio-Visual Fusion. The temporal fusion branch models relationships between audio and visual content across time. We apply two convolutional layers to integrate the embedding from each modality at each time step into a single token. The resulting visual and audio tokens are denoted as $z_{v,t}\in\mathbb{R}^{T\times 1\times D}$ and $z_{a,t}\in\mathbb{R}^{T\times 1\times D}$, respectively. Then we feed $z_{t}=\left[z_{v,t},z_{a,t}\right]\in\mathbb{R}^{2T\times 1\times D}$ into a cross-frame self-attention layer $\pi$ that can be formulated as:
$\pi(z_{t})=Softmax\left(\bm{Q}_{t}\bm{K}_{t}^{T}/\sqrt{D}\right)\bm{V}_{t}\in\mathbb{R}^{2T\times 1\times D},$ (6)
@@
Merging of Two Fusion Modules. After obtaining audio-visual representations from the two fusion modules, we merge the two branches by reweighting the output from spatial fusion with the temporal weights from temporal fusion in each channel, which produces a new representation for each modality that has been refined by multimodal spatial and temporal correlation. Specifically, we break down the output from spatial fusion $u_{s}\in\mathbb{R}^{T\times(N+1)\times D}$ into $u_{v,s}\in\mathbb{R}^{T\times N\times D}$ and $u_{a,s}\in\mathbb{R}^{T\times 1\times D}$, and the output from temporal fusion $u_{t}\in\mathbb{R}^{2T\times 1\times D}$ into $u_{v,t}\in\mathbb{R}^{T\times 1\times D}$ and $u_{a,t}\in\mathbb{R}^{T\times 1\times D}$. The reweighted visual representation is formulated as
$\vspace{-0.04cm}u_{v}=u_{v,s}\otimes u_{v,t}\in\mathbb{R}^{T\times N\times D},\vspace{-0.04cm}$ (7)